OpenStudy (anonymous):

What rule changes the input numbers to output numbers? Input - 1, 2, 3, 4 Output- -2, 1, 4, 7 A. x-3 B. 2x-4 C. 3x - 5 D. 4x - 6

OpenStudy (joemc):

The key to answering a question like this is first, understanding what you are being asked. Then, approaching it in a logical manner. You are given a list of input, and output numbers. Then equations. You need to substitute each input number in for the 'x' in each equation. The correct answer will give you the corresponding output number for each input number in the question. So, start with choice 'A', x-3 Your first input number is 1. Substitute this in for the 'x'. So, x-3 becomes 1-3. The result is -2. That is the output value for 1. So far, so good. Go on to the next input value: 2... Substitute this in for 'x'. x-3 becomes 2-3, which is -1. This is a problem, the output value for 2 should be 1. So, that eliminates x-3. Go on to the next possible answer: 2x-4. Let's try this with our first input, 1. 2x-4 becomes 2(1) -4 = 2-4 = -2. So far, so good. The next input value given is 2. 2x-4 becomes 2(2)-4, or 4-4=0. That is not equal to the output value for 2, which is 1. So, that eliminate that possible answer. Keep working this way through the two other choices given. One of the remaining choices will give you the output value given for each of the four input values.